# Optimal smoothing parameter in Eilers-Wittaker smoother

### v1

@@ I Introduction
Eilers introduced a smoothing procedure based on a linear optimization procedure with regularisation eilers03 . In the so-called Whittaker–Eilers method, two terms are used in a minimization procedure, one takes care of reducing the mean square error between the smoothed data and the original signal. In contrast, the second regularisation term is a penalty term that intends to smooth the solution, effectively reducing noise at the expense of the first term. A regularisation parameter governs the balance between the two terms. The selection of the optimal regularisation parameter remains a critical issue: excessive values may suppress essential features, while insufficient values fail to effectively attenuate noise.
@@
II Whittaker-Eilers algorithm
The starting point is noisy data $y(t)$, which is the experimental realization of an underlying smooth signal $s(t)$ corrupted with noise as described by equation ()1). The residual of the experimental data and the underlying smooth data will be given by
$R=\sum\limits_{t}(y(t)-\hat{s}(t))^{2},$ (2)
whereas $\hat{s}(t)$ is an approximation of $s(t)$ upon smoothing. Smoothing can be measured by the n-order numerical derivative of the experimental data. Up to the first order, this will be given by
$S_{1}=\sum\limits_{q}(\hat{s}(t)-\hat{s}(t-1))^{2},$ (3)
@@ II Whittaker-Eilers algorithm
where, for first order differences,
$D=\left(\begin{array}[]{ccccccc}-1&1&0&0&\ldots&0&0\\ 0&-1&1&0&\ldots&0&0\\ 0&0&-1&1&\ldots&0&0\\ &&\cdot&\cdot&\cdot&&\\ 0&0&0&&\ldots&-1&1\end{array}\right).$ (7)
@@ III The regularization parameter
The $\lambda$ parameter in (5) determines the smoothing behaviour of the Eilers algorithm and therefore, how to choose it becomes an important step in the smoothing procedure. The goal of selection procedures is to provide a systematic method for choosing the smoothing parameter, $\lambda$, rather than relying on subjective visual assessment. Several procedures are used for this selection, and, for completeness, we describe two of the most commonly used ones.
@@ IV Spectral entropy selection criterion
The corresponding $e_{\lambda}$ vs $\lambda$ curve will be called the S-curve. In Figure 2, the S-curve of an arbitrary continuous function is shown. The curve shows local maximum values. The $\lambda$ value for the absolute maximum of $e_{\lambda}$ is taken as the regularization parameter.
@@ V.1 Experimental data
Figure 5: Experimental Datasets. Application and effectiveness of the proposed spectral entropy method on three distinct types of real-world experimental data. Upper Panel (Sugar Stocks): The data represents a time series of sugar stock prices from the commodities market. Middle Panel (Galaxy Data): This plot shows optical single-fiber spectroscopy data from the Sloan Digital Sky Survey. Lower Panel (NMR Data): The data is from a Nuclear Magnetic Resonance (NMR) experiment. Although the Whittaker-Eilers smoother is known to sometimes underestimate the intensity of very narrow peaks, the $\lambda$ selected by the spectral entropy method allows the smoothed curve to correctly identify the position of each peak’s maximum. In each panel, the right inset is the S-curve obtained for its corresponsind dataset. The examples serve to validate the proposed method, showing that it produces well-behaved and visually compatible curves for data from different domains.
@@
Figure 6: Difference Operator Order. The impact of the choice of order for the difference matrix $D$ in the behavior of the Whittaker-Eilers smoother, with the regularization parameter $\lambda$ in each case selected by the proposed spectral entropy method. The analysis is performed on a time series of Euro to Dollar conversion rates, a dataset characterized by abrupt changes and an overall increasing trend. Order 1 (Left Panel): The first-order difference operator produces a smoothed curve that captures the general, overall trend of the data. However, it is not sensitive enough to follow the local variations, missing the smaller peaks and troughs present in the time series. Order 2 (Middle Panel): Using a second-order difference operator results in a significant improvement. The smoothed curve is now able to follow all the local maxima and minima, effectively capturing the detailed fluctuations in the data while still providing a smooth and visually pleasing result. Order 3 (Right Panel): The third-order smoother behaves similarly to the second-order one. While it follows the local variations slightly more closely, the overall change from order 2 is not substantial. It continues to strike a good balance, capturing important details without excessively overfitting the noise in the underlying data.
@@ VII Acknowledgments
CITMA partially funded this work under the project CARDENT (PN223LH010-053).

### v2

# Optimal smoothing parameter in Eilers-Whittaker smoother

@@ I Introduction
Eilers introduced a smoothing procedure based on a linear optimization procedure with regularisation eilers03 . In the so-called Eilers-Whittaker method, two terms are used in a minimization procedure, one takes care of reducing the mean square error between the smoothed data and the original signal. In contrast, the second regularisation term is a penalty term that intends to smooth the solution, effectively reducing noise at the expense of the first term. A regularisation parameter governs the balance between the two terms. The selection of the optimal regularisation parameter remains a critical issue: excessive values may suppress essential features, while insufficient values fail to effectively attenuate noise.
@@
II Eilers-Whittaker algorithm
The starting point is noisy data $y(t)$, which is the experimental realization of an underlying smooth signal $s(t)$ corrupted with noise as described by equation (1). The residual of the experimental data and the underlying smooth data will be given by
$R=\sum\limits_{t}^{n}(y(t)-\hat{s}(t))^{2},$ (2)
whereas $\hat{s}(t)$ is an approximation of $s(t)$ upon smoothing. $n$ is the number of experimental sampled points. Smoothing can be measured by the n-order numerical derivative of the experimental data. Up to the first order, this will be given by
$S_{1}=\sum\limits_{q}^{n}(\hat{s}(t)-\hat{s}(t-1))^{2},$ (3)
@@ II Eilers-Whittaker algorithm
where, for first order differences, $D$ is a $(n-1)\times n$ matrix,
$D=\left(\begin{array}[]{ccccccc}-1&1&0&0&\ldots&0&0\\ 0&-1&1&0&\ldots&0&0\\ 0&0&-1&1&\ldots&0&0\\ &&\cdot&\cdot&\cdot&&\\ 0&0&0&0&\ldots&-1&1\end{array}\right),$ (7)
and for second order, the $(n-2)\times n$ matrix,
$D=\left(\begin{array}[]{ccccccccc}1&-2&1&0&0&\ldots&0&0&0\\ 0&1&-2&1&0&\ldots&0&0&0\\ 0&0&0&1&-2&\ldots&0&0&0\\ &&&\cdot&\cdot&\cdot&&\\ 0&0&0&0&0&\ldots&1&-2&1\end{array}\right),$ (8)
@@ III The regularization parameter
The $\lambda$ parameter in expression (5) determines the smoothing behaviour of the Eilers algorithm and therefore, how to choose it becomes an important step in the smoothing procedure. The goal of selection procedures is to provide a systematic method for choosing the smoothing parameter, $\lambda$, rather than relying on subjective visual assessment. Several procedures are used for this selection, and, for completeness, we describe two of the most commonly used ones.
@@ IV Spectral entropy selection criterion
The corresponding $e_{\lambda}$ vs $\lambda$ curve will be called the S-curve. In Figure 2, the S-curve of an arbitrary continuous function is shown. The curve shows local maximum values. The $\lambda$ value for the absolute maximum of $e_{\lambda}$ was empirically found to be a good choice for the regularization parameter.
@@ V.1 Experimental data
Figure 5: Experimental Datasets. Application and effectiveness of the proposed spectral entropy method on three distinct types of real-world experimental data. Upper Panel (Sugar Stocks): The data represents a time series of sugar stock prices from the commodities market. Middle Panel (Galaxy Data): This plot shows optical single-fiber spectroscopy data from the Sloan Digital Sky Survey. Lower Panel (NMR Data): The data is from a Nuclear Magnetic Resonance (NMR) experiment. Although the Eilers-Whittaker smoother is known to sometimes underestimate the intensity of very narrow peaks, the $\lambda$ selected by the spectral entropy method allows the smoothed curve to correctly identify the position of each peak’s maximum. In each panel, the right inset is the S-curve obtained for its corresponsind dataset. The examples serve to validate the proposed method, showing that it produces well-behaved and visually compatible curves for data from different domains.
@@
Figure 6: Difference Operator Order. The impact of the choice of order for the difference matrix $D$ in the behavior of the Eilers-Whittaker smoother, with the regularization parameter $\lambda$ in each case selected by the proposed spectral entropy method. The analysis is performed on a time series of Euro to Dollar conversion rates, a dataset characterized by abrupt changes and an overall increasing trend. Order 1 (Left Panel): The first-order difference operator produces a smoothed curve that captures the general, overall trend of the data. However, it is not sensitive enough to follow the local variations, missing the smaller peaks and troughs present in the time series. Order 2 (Middle Panel): Using a second-order difference operator results in a significant improvement. The smoothed curve is now able to follow all the local maxima and minima, effectively capturing the detailed fluctuations in the data while still providing a smooth and visually pleasing result. Order 3 (Right Panel): The third-order smoother behaves similarly to the second-order one. While it follows the local variations slightly more closely, the overall change from order 2 is not substantial. It continues to strike a good balance, capturing important details without excessively overfitting the noise in the underlying data.